Determine the area of the region bounded by the given set of curves y=x^2+2, y=sin(x), x=2.


1
Expert's answer
2021-04-19T05:04:04-0400

"Assuming\ it\ is\ also\ bounded\ by\ x=0,\"

"\int_{0}^{2}{x^2+2-\sin{\left(x\right)}\ dx}"

"\frac{x^3}{3}+2x+\cos{\left(x\right)}\ \ \left(0,\ 2\right)"

"\frac{8}{3}+4+\cos{\left(2^c\right)}-1"

"5.2505\ square\ units"